OpenStudy (anonymous):

If 3x^2-4x+7=0, then (x-(2/3)^2 = a) 17/9 b) -17/9 c) 5/9 d) -5/9 answer is b Please explain how to get to this answer, thank you

OpenStudy (anonymous):

\(3x^2 -4x +7=0\) divide both sides by 3, you have \(x^2−\dfrac{4}{3}x+\dfrac{7}{3}=0\) It can be written as \( x^2−2*\dfrac{2}{3}x+\dfrac{7}{3}=0\) Ok so far?

OpenStudy (anonymous):

not finish yet, but if you don't get that part, I can't do more.:(